Calculate the angles, side lengths and area of the triangle with vertices at the points A = (1, 2, 1), B = (4, 8, 3), C = (7, 0, −2).
Answer #1

Sol Giren At A(1,2,1). B14,8,3). C170,-2) Distance formule TPB! = 162,-X1] + 19-4,32+ (37-7,52 7/ . \7 LABI = (4-11²+(2-8)+(3